# Towards Data-Driven Metrics for Social Robot Navigation Benchmarking

### v1

@@ Abstract
This paper presents a joint effort towards the development of a data-driven Social Robot Navigation metric to facilitate benchmarking and policy optimization. We provide our motivations for our approach and describe our proposal for storing rated social navigation trajectory datasets. Following these guidelines, we compiled a dataset with 4427 trajectories—182 real and 4245 simulated—and presented it to human raters, yielding a total of 4402 rated trajectories after data quality assurance. We also trained an RNN-based baseline metric on the dataset and present quantitative and qualitative results. All data, software, and model weights are publicly available.
@@ I Introduction
Social robot navigation is becoming an increasingly relevant field, with both growing academic research interest and notable industry involvement. Industrial applications include autonomous vehicles, robots for restaurants and museums, and delivery pods for last mile delivery [15].
Figure 1: We acquire trajectories from both real and simulated scenarios. Simulated trajectories have a lower cost and allow for ethical recording of unsafe behavior, but real trajectories are still required to ensure generalisation. We generate a top-down view for each trajectory and show it to raters to collect scores. We have made all the data we collected public, including trajectories, raters’ data, and their ratings. Metric and policy learning are use cases for the dataset.
@@
Despite efforts of the social robot navigation community to establish principles and guidelines for evaluation [4], benchmarking of social robotics is not yet standardized and can be biased to suit specific research narratives. Social robot navigation (hereafter SocNav) is currently assessed through metrics focusing on success, safety, proxemics, or readability [10]. However, different research works consider different metrics depending on what their authors consider relevant and often make claims without considering tradeoffs between metrics or Pareto optimality for the metrics selected.
@@
Furthermore, task type and navigation context [4] are often overlooked. How the type of task affects behavior is exemplified by robots entering people’s personal space, which is required if a robot is delivering an object to a person, but normally should be avoided. The importance of context is exemplified by whether a robot should pass bystanders at high speed, which is more appropriate when a robot is carrying urgent medical equipment than when returning to base to recharge, even though the task might be the same from a geometric perspective.
@@
Many existing metrics are computed algorithmically from variables, or analytic [4]. Analytic metrics often have poor variable-wise scalability, and disregard factors such as the type of task or contextual factors such as the distribution of objects around the robot. Analytic metrics also frequently lack experimental support showing that what the metric measures matches user perceptions of robot behavior quality. Nevertheless, new metrics continue to be proposed, and, although the intuitions behind them are generally solid, they also often lack empirical support. We argue that metrics should be empirically validated, ideally learned from data, and should evaluate the overall effectiveness and socialness of an entire trajectory. We call such a metric an All-encompassing Learned Trajectory-wise (ALT) metric [4]. An ALT metric is learned directly from human evaluations of full robot trajectories, conditioned on context—a category of metrics which, to our knowledge, does not yet exist.
@@
Motivated by the lack of a standardized method to benchmark SocNav policies, limitations of existing analytic metrics, and the absence of an ALT metric, this article studies the feasibility of a data-driven solution to create an ALT metric, which would solve all these problems.
@@
Robots are not yet commonly integrated into everyday human environments with explicit quantitative feedback [20], so few datasets suitable for training learned metrics exist. This gap makes compiling a dataset specifically for this purpose essential. Therefore, this paper contributes with:
@@
Our proof-of-concept metric shows that an ALT metric can be learned, and encourages the community to help scale the dataset to the point where metrics learned over it can be used as the de facto metric for benchmarking robot trajectories [^1] Instructions to contribute are in the software repository https://github.com/SocNavData/SocNav3.. Policy development is out of the scope of this paper.
In addition to our main contributions, we provide experimental results on data quality and the accuracy of our metric, as well as qualitative results for stereotypical scenarios.
@@ II Related Work
Regarding learned SocNav metrics, SocNav1 [8] and SocNav2 [2] are the initiatives that most closely resemble the goals of the paper at hand. The SocNav1 [8] dataset provides scores of static scenarios including one robot and a variable number of humans, with generic object bounding boxes and explicit interactions (human-to-human and human-to-object). The SocNav2 dataset [2] extended SocNav1 to account for brief periods of time and specific goals, but not full trajectories. Even though these works provide learned SocNav metrics, they are step-wise, i.e., they account for snapshots rather than whole trajectories. Furthermore, none of them consider the context of the tasks.
Although not strictly a metric, another notable research direction estimates the perturbation of pedestrian movement as a variable to minimize  [5, 1]. Despite its elegant simplicity, this approach is dependent on the number of people in the environment, and does not account for the context of the task nor the general perception of trajectory quality. Similarly, [13] learns a cost function over observations of real human motion interactions found in publicly available datasets, then uses that cost to determine a navigation controller. While this cost function could be used as a metric, this approach only considers the context of crossing pedestrians on a pavement and has the same shortcomings as [5] and [1].
Many datasets relevant to human trajectory modeling exist, but datasets which also include humans, robots and environmental data (such as a map or objects) are scarcer. JRDB [9], THOR [14], SCAND [7] and SACSoN [5] are examples of such datasets. They provide raw sensor data from teleoperated or autonomous robot trajectories, and frequently include a video feed and 3D LiDAR or RGBD data. However, with the exception of THOR [14], pedestrians are annotated as bounding boxes or 2D points without orientation.
We found no datasets containing rated robot trajectories, so we chose for this paper to reuse and annotate existing datasets along with recording new additional data. Although THOR initially seemed a good fit for this purpose because it includes object annotations, oriented pedestrians and even human gaze orientation, we chose SACSON instead because of its scale. Nevertheless, as mentioned in the introduction, we combined synthetic data with this real data so we could vary the quality of the trajectories in order to reduce the bias of the learned metric.
@@ III Dataset Specification and Tools
This section first describes our proposed dataset specification, including the variables it represents and its file format, and then summarizes the software we developed to work with compatible datasets. The design decisions behind this specification are the result of a series of focus group sessions where the authors of the paper participated.
@@ III-A Variables
The dataset comprises variables related to raters, trajectories, and rater-trajectory scores. For every rater, the dataset contains their age (integer), gender (categorical: female, male, non-binary, transgender, other, no-answer), country of residence (categorical), and a rating list. Each rating list contains tuples $(t,c,r)$, where $t$ is a trajectory identifier (string), $c$ is a context (string), and a $r$ is a score (float, $r\in[0,1]\subset\mathbb{R}$). We chose to acquire real-valued scores rather than a Likert scale following best-practice HRI studies [18] which regard sensitivity as more relevant than repeatability. To score trajectories, the survey tool used a slider with three references “extremely bad” (score 0), fair (score $0.5$) and “extremely good” (score $1$). Note that contexts are not necessarily bound to trajectories when trajectories are recorded; this enables us to vary the context to explore how different contexts affect the perception of the same trajectory.
@@
The robot is described providing its pose (in $SE(2)$), speed (a twist vector), drive system (categorical variable, “differential”, “omnidirectional”, “ackerman”, or “biomimetic”) and shape (circle, rectangle, or a generic polygon described using a vertex representation). The focus group considered whether to add 3D information about the shape of robots or objects. While this could help detect potentially problematic situations (e.g., robot or object geometry impeding the visibility of other agents), most current metrics are very far from considering such aspects. Therefore, the focus group decided not to include 3D shapes in the initial version of the specification.
The goal is defined differently depending on the task type, which is provided as a categorical variable “go-to”, “guide-to”, “follow”, or “interact-with”. For go-to tasks, the specification requires a target 2D position (in meters) and a target angle (in radians), both with thresholds using Euclidean and angular distance, respectively. By varying these threshold values, a go-to task can specify a very stringent $SE(2)$ point target, just a position, or only an angle. Tasks of type guide-to, are specified similarly to go-to tasks, but require an additional human identifier, and the task target applies to the human, not the robot. Tasks of type follow and interact-with require a human identifier to specify which human is the robot expected to follow, along with an Euclidean distance threshold.
@@
Type Task type, either “go-to", “guide-to" or “follow".
@@
Environment Area semantics Free text describing the area -+ e.g., “indoor", “outdoor", “kitchen", “a science museum".
@@ III-B Structure and storage
The raw dataset contains two main directories: one containing the trajectories and another one containing data about the raters and their ratings.
The trajectories directory contains JSON files of recorded trajectories in sub-directories named according to the source of the trajectory data. Each file contains one trajectory, and their relative file paths serve as trajectory identifiers. The data structure of these JSON files is presented in Table I.
The ratings directory contains a separate JSON file for each questionnaire completed, as described in III-A. In addition to the rater’s demographic information, each of these files store their ratings as a list of tuples: trajectory identifier (i.e., a relative path to the JSON file), context, score.
@@ III-C Software layer
Trajectory visualization tool: Unlike other types of datasets with simpler datatypes, such as plain text, images, or audio, trajectory data combines different variables, which makes it harder to work with and visualize. Visualization is crucial in this domain because the complexity of the data structures makes errors likely if the trajectories are not carefully compiled. Inconsistencies in coordinate or unit systems are common issues that can be picked up visually early on with the right tools. The visualization tool is shown in 2(a). A 3D visualization tool built on top of Webots [11] is also provided to generate video recordings.
(a) 2D custom representation.
(b) Webots-based 3D representation.
Figure 2: Trajectory visualization using a custom 2D representation (2(a)) and 3D simulator (Webots 2(b)).
@@
Surveying tool: Running a ratings survey benefits from a more usable—and aesthetic—representation. The surveying tool shows video recordings along with a randomly generated context, asks raters to provide their scores, as shown in Fig. 3. The trajectories to use, the types of contexts, the maximum number of scores per rater, and the repeated control questions can all be configured.
Figure 3: Survey screenshot.
Data normalization, and augmentation: Normalization and augmentation help to include equivariances and invariances in machine learning models without embedding them into equations. Applying these transformations to robot trajectories is complex and error-prone, and so, as with visualization, tooling support is helpful. To avoid forcing users to duplicate data augmentation code, and to streamline and homogenize data augmentation and normalization, PyTorch data transforms implementing such features are provided.
Data normalization automatically reframes all poses into the goal frame of reference; this can steer machine learning models towards invariance w.r.t. the frame of reference.
@@
The provided software implements three types of data augmentations. First, to enhance robustness to noise, poses can be augmented with additive Gaussian noise. Secondly, when goal orientations exceed a threshold larger or equal to $\pi$ rad, the orientation value is randomized. Finally, trajectory data can be randomly mirrored (left-right).
IV Description and Analysis of the Dataset Prototype
@@
We provide a dataset to show how the proposed dataset structure can be exploited to generate a learning-based metric. We collected $4427$ trajectories, $182$ real and $4245$ simulated. We extended this initial set by varying the configuration of the walls in the scenarios of the compiled trajectories (i.e., removing some or all of the walls). This process resulted in a total of $7472$ trajectories. It is key to have both real and simulated data sources because they pose different challenges: real trajectories typically have lower variability than simulated ones because of the expense of recording trajectories in multiple scenarios, while simulated data have less realistic movements and noise. Additionally, due to safety and ethical reasons, dangerous robot trajectories can only be acquired in simulation. Furthermore, acquiring a large enough dataset to cover all possible situations with real data exclusively would be unrealistically expensive.
@@
In this study, we kept cost bounded by limiting this proof-of-concept dataset in several ways, nevertheless ensuring good coverage within the selected domain. Firstly, the dataset is limited to indoor areas whose bounding box ranges between 6m6~m and 10m10~m wide. The navigation tasks covered in the dataset are all of the type “go-to”; other tasks such as following or guiding humans are left for future additions to the dataset repository. The objects considered in the dataset are limited to chairs, shelves, tables, lights, and plants. The maximum robot speed recorded is 2m/s2~m/s.
@@ IV-A Trajectories recorded
Two sources were used to acquire real trajectories. A set of 109 trajectories was recorded at ARP laboratory (Aston University). A $8\times 7.4$ meters wide room was populated with tables, shelves, chairs, a TV, and from 1 to 6 humans. Another 73 trajectories were re-used from the SACSoN dataset [5]. The trajectories available in SACSoN were recorded in environments larger than the maximum area set for the proof-of-concept dataset, so they were cropped to comply with the range allowed within the restrictions for our dataset.
We included simulated trajectories from a range of simulators and algorithms, using both manual control and a learned RL agent on SocNavGym [6], HATEB [16], and HuNavSim [12], as well as manual control on Gazebo.
@@ IV-B Data quality
Each survey included $15$ control questions to assess consistency between the participants. Additionally, five of these questions were presented twice to evaluate intra-rater consistency. In total, $49$ participants scored $6,481$ trajectories.
To ensure data quality, we first discarded ratings from participants who did not answer all control questions. For the $34$ remaining participants, we then applied a selection process that considers both intra-rater and inter-rater consistency using the quadratic weighted Cohen’s kappa coefficient. We define $\mu_{intraH}$ as a high-quality threshold and $\mu_{intraL}$ as a minimum-quality threshold for intra-rater consistency and $\mu_{inter}$ as a minimum threshold for inter-rater consistency. The selection consists of the following steps:
@@
    Discard raters with an intra-rater consistency lower than $\mu_{intraL}$ or an inter-rater consistency lower than $\mu_{inter}$ with respect to the average scores of the control questions.
Combining intra- and inter-rater agreements provides more solid results than a separate selection. Due to the limited number of repeated control questions, small variations in the answers of a rater could have a significant impact on their intra-rater agreement [19]. Thus, an initial selection based only on high intra-rater consistency could result in discarding too many raters. Using both coefficients, we can retain raters who align with the group opinion, even with minor inconsistencies in their responses. This approach improves the overall reliability and representativeness of the data by filtering out erratic and atypical ratings without being excessively restrictive.
Figure 4 shows the consistency map of the raters selected following the described procedure, considering values of $0.4$, $0.1$ and $0.2$ for $\mu_{intraH}$, $\mu_{intraL}$ and $\mu_{inter}$, respectively. The resulting set of 22 raters yields a final dataset of 4402 scored trajectories.

### v2

@@ Abstract
This paper presents a joint effort towards the development of a data-driven Social Robot Navigation metric to facilitate benchmarking and policy optimization for ground robots. We compiled a dataset with 4427 trajectories —182 real and 4245 simulated— and presented it to human raters, yielding a total of 4402 rated trajectories after data quality assurance. Notably, we provide the first all-encompassing learned social robot navigation metric, along qualitative and quantitative results, including the test loss achieved, a comparison against hand-crafted metrics, and an ablation study. All data, software, and model weights are publicly available.
@@ I Introduction
Social robot navigation is an increasingly relevant field, with both growing academic research interest and notable industry involvement. Despite the efforts of the social robot navigation community to establish principles and guidelines for evaluation [francis2023principles], benchmarking of social robots is not yet standardized and can be biased to suit specific research narratives. Social robot navigation (hereafter SocNav) is currently assessed through metrics focusing on success, safety, proxemics, or readability [mavrogiannis2023core]. However, different research works use different metrics depending on what their authors consider relevant and often make claims without contemplating tradeoffs between metrics or Pareto optimality.
Figure 1: We acquire trajectories from both real and simulated scenarios. Simulated trajectories have a lower cost and allow for ethical recording of unsafe behavior, but real trajectories are still required to ensure generalization. We generate a top-down view for each trajectory and show it to raters to collect scores. All the data collected is public, including trajectories, raters’ data, and their ratings. Metric and policy learning are use cases for the dataset.
@@
Furthermore, task type and navigation context [francis2023principles] are often overlooked. How the type of task affects behavior is exemplified by robots entering people’s personal space, which is required if a robot is delivering an object to a person, but normally should be avoided. The importance of context is exemplified by whether a robot should pass bystanders at high speed, which is more appropriate when a robot is carrying urgent medical equipment than when returning to base to recharge, even though the task might be the same from a geometric perspective.
@@
Many existing metrics are analytic, computed algorithmically from variables, which leads to poor variable-wise scalability, and disregard factors such as the type of task or contextual factors such as the distribution of objects around the robot [francis2023principles]. Furthermore, they frequently lack experimental support showing that what the metric measures matches user perceptions of robot behavior quality. Nevertheless, new metrics continue to be proposed, and, although the intuitions behind them are generally solid, they also often lack empirical support. We argue that metrics should be empirically validated, ideally learned from data, and should evaluate the overall effectiveness and socialness of an entire trajectory. We call such a metric an All-encompassing Learned Trajectory-wise (ALT) metric [francis2023principles]. An ALT metric is learned directly from human evaluations of full robot trajectories, conditioned on context —a category of metrics which, to our knowledge, does not yet exist.
@@
Motivated by the lack of a standardized method to benchmark SocNav policies, limitations of existing analytic metrics, and the absence of an ALT metric, this article studies the feasibility of a data-driven solution to create an ALT metric, which would solve all the aforementioned problems.
@@
Robots are not yet commonly integrated into everyday human environments with explicit quantitative feedback [zhang2025predicting], so no datasets suitable for training learned metrics exist. This gap makes compiling a dataset specifically for this purpose essential. Therefore, this paper contributes with:
@@
Our proof-of-concept metric shows that an ALT metric can be learned, and encourages the community to help scale the dataset to the point where metrics learned with it can be used as the de facto metric for benchmarking robot trajectories [^1] Instructions to contribute are in the software repository https://github.com/SocNavData/SocNav3. Policy development is out of the scope of this paper.
The experimental results in Sec. V provide evidence of the quality of the data compiled and the metric developed. They include the test loss achieved, an ablation study, an analysis of the correlation between our metric and some of the most common SocNav metrics, and qualitative results for stereotypical scenarios.
@@ II Related Work
Regarding learned SocNav metrics, SocNav1 [manso2020socnav1] and SocNav2 [socnav2_multimedia2022] are the initiatives that most closely resemble the goals of this paper. SocNav1 [manso2020socnav1] provides scores of static scenarios including one robot and a variable number of humans, with generic object bounding boxes and explicit interactions (human-to-human and human-to-object). SocNav2 [socnav2_multimedia2022] extended SocNav1 to account for brief periods of time and specific goals, but not full trajectories. Although these works provide learned SocNav metrics, they do not consider the context of the tasks, and are step-wise, i.e., they account for snapshots rather than whole trajectories.
Although not strictly a metric, another notable research direction estimates the perturbation of pedestrian movement as a variable to minimize  [hirose2023sacson, agrawal2024evaluating]. Despite its elegant simplicity, this approach is dependent on the number of people in the environment, and does not account for the context of the task nor the general perception of trajectory quality. Similarly, [perez2014robot] learns a cost function over observations of real human motion interactions found in publicly available datasets, then uses that cost to determine a navigation controller. While this cost function could be used as a metric, this approach only considers the context of crossing pedestrians on a pavement and has the same shortcomings as [hirose2023sacson] and [agrawal2024evaluating].
Many datasets relevant to human trajectory modeling exist, but datasets which also include humans, robots and environmental data (such as a map or objects) are scarce. JRDB [martin2019jrdb], THOR [thorDataset2019], SCAND [scanddataset] and SACSoN [hirose2023sacson] are examples of such datasets. They provide raw sensor data from teleoperated or autonomous robot trajectories, and frequently include a video feed and 3D LiDAR or RGBD data. However, with the exception of THOR [thorDataset2019], pedestrians are annotated as bounding boxes or 2D points without orientation. Orientation can be very informative as it captures interaction intents [singamaneni2023survey], even when humans are static.
We found no datasets containing rated robot trajectories, so we chose to reuse and annotate existing datasets along with recording additional data. Although THOR includes over 60 minutes of recordings including object annotations, oriented pedestrians and human gaze orientation, we chose to integrate SACSON first because of its scale (75 hours of recordings) and variability (i.e., the trajectories in THOR were all recorded in the same space. As mentioned in the introduction, we combine synthetic data with real data to vary the quality of the trajectories in order to reduce the bias of the learned metric towards high scores. In this vein, it must be noted that the new trajectories that we recorded include robot behavior of lower quality than those in existing datasets, contributing to reducing the bias toward higher scores.
@@ III Dataset Specification and Tools
This section first describes our proposed dataset specification, including the variables it represents and its file format, and then summarizes the software we developed to work with compatible datasets.
@@ III-A Variables
The dataset comprises variables related to raters, trajectories, and rater-trajectory scores. For every rater, the dataset contains their age (integer), gender (categorical: female, male, non-binary, transgender, other, no-answer), country of residence (categorical), and a rating list. Each rating list contains tuples $(t,c,r)$, where $t$ is a trajectory identifier (string), $c$ is a context (string), and a $r$ is a score (float, $r\in[0,1]\subset\mathbb{R}$). We chose to acquire real-valued scores rather than a Likert scale following best-practice HRI studies [taylor2022observer] which regard sensitivity as more relevant than repeatability. To score trajectories, the survey tool used a slider with three references “extremely bad” (score 0), fair (score $0.5$) and “extremely good” (score $1$). Note that contexts are not necessarily bound to trajectories when trajectories are recorded; this enables us to vary the context to explore how different contexts affect the perception of the same trajectory.
@@
The robot is described providing its pose (in $SE(2)$), speed (a twist vector), drive system (categorical variable, “differential”, “omnidirectional”, “ackerman”, or “biomimetic”) and shape (circle, rectangle, or a generic polygon described using a vertex representation).
The goal is specified depending on the task type, which is a categorical variable (“go-to”, “guide-to”, “follow”, or “interact-with”). For go-to tasks, the specification requires a target 2D position (in meters) and a target angle (in radians), both with Euclidean and angular distance thresholds, respectively. By varying threshold values, a go-to task can specify a very stringent $SE(2)$ point target, just a position, or only an angle. Tasks of type guide-to, are specified similarly to go-to tasks, but require an additional human identifier, and the task target applies to the human, not the robot. Tasks of type follow and interact-with require a human identifier to specify which human is the robot expected to follow, along with an Euclidean distance threshold.
@@
Type Task type, either “go-to”, “guide-to” or “follow”.
@@
Environment Area semantics Free text describing the area e.g., “indoor”, “outdoor”, “kitchen”, “a science museum”.
@@ III-B Structure and storage
The raw dataset contains two main directories: one containing the trajectories and another one containing data about the raters and their ratings. The trajectories directory contains JSON files of recorded trajectories in sub-directories named according to the source of the trajectory data. Each file contains one trajectory, and their relative file paths serve as trajectory identifiers. The data structure of these JSON files is presented in Table I. The ratings directory contains a separate JSON file for each questionnaire completed, as described in III-A. In addition to the rater’s demographic information, each of these files store their ratings as a list of tuples: trajectory identifier (i.e., a relative path to the JSON file), context, score.
@@ III-C Software layer
Trajectory visualization tool: Unlike other types of datasets with simpler datatypes, such as images, or audio, SocNav data combines different variables, which makes errors likely if the trajectories are not carefully compiled. Inconsistencies in coordinate or unit systems are common issues that can be picked up with the tools provided.
@@
Surveying tool: Running a ratings survey benefits from a more usable —and aesthetic— representation. The surveying tool shows video recordings along with randomly generated contexts, and asks raters to provide their scores.
@@
Data normalization, and augmentation: Normalization and augmentation help to include equivariances and invariances in machine learning models without embedding them into equations. To avoid forcing users to duplicate complex and error-prone data augmentation code, and to streamline and homogenize data augmentation and normalization, data transforms are provided. Data normalization automatically reframes all poses into the goal frame of reference; this can steer machine learning models towards invariance w.r.t. the frame of reference. The provided software implements three types of data augmentations. First, to enhance robustness to noise, poses can be augmented with additive Gaussian noise. Secondly, when goal orientations exceed a threshold larger or equal to $\pi$ rad, the orientation value is randomized. Finally, trajectory data can be randomly mirrored (left-right).
IV Description and Analysis of the Dataset
@@
We provide a dataset to show how the proposed dataset structure can be exploited to generate a learned metric. We collected $4427$ trajectories, $182$ real and $4245$ simulated. We extended this initial set by varying the configuration of the walls in the scenarios of the compiled trajectories (i.e., removing some or all of the walls). This process resulted in a total of $7472$ trajectories. It is key to have both real and simulated data sources because they pose different challenges: real trajectories typically have lower variability than simulated ones because of the expense of recording trajectories in multiple scenarios, while simulated data have less realistic movements and noise. Additionally, due to safety and ethical reasons, dangerous robot trajectories can only be acquired in simulation. Furthermore, acquiring a large enough dataset to cover all possible situations with real data exclusively would be unrealistically expensive.
@@
Given resource constraints and the goal of focusing on practically relevant scenarios with good coverage, we implemented focused sampling. We selected the domain of indoor areas whose bounding box sides range between 6m6~m and 10m10~m. The navigation tasks covered in the dataset are all of the type “go-to”; other tasks such as following or guiding humans are left for future additions to the dataset repository. The objects considered in the dataset are limited to chairs, shelves, tables, lights, and plants. The maximum robot speed recorded is 2m/s2~m/s.
@@ IV-A Trajectories recorded
Two sources were used to acquire real trajectories. A set of 109 trajectories was recorded at ARP laboratory (Aston University). A $8\times 7.4$ meters wide room was populated with tables, shelves, chairs, a TV, and from 1 to 6 humans. Another 73 trajectories were re-used from the SACSoN dataset [hirose2023sacson]. The trajectories available in SACSoN were recorded in environments larger than the maximum area set for the dataset, so they were cropped to comply with the range allowed within the restrictions for the dataset.
We included simulated trajectories from a range of simulators and algorithms, using both manual control and a learned RL agent on SocNavGym [kapoor2023socnavgym], HATEB [singamaneni2021human], and HuNavSim [perez2023hunavsim], as well as manual control on Gazebo.
@@ IV-B Data quality
Each survey included $15$ control questions to assess consistency between the participants. Additionally, five of these questions were presented twice to evaluate intra-rater consistency. In total, $49$ participants scored $6481$ trajectories.
To ensure data quality, we first discarded ratings from participants who did not answer all control questions, as it would not be possible to analyze their consistency. For the $34$ remaining participants, we applied a selection process that considers both intra-rater and inter-rater consistency using the quadratic weighted Cohen’s kappa coefficient. We define $\mu_{intraH}$ as a high-quality threshold and $\mu_{intraL}$ as a minimum-quality threshold for intra-rater consistency and $\mu_{inter}$ as a minimum threshold for inter-rater consistency. The selection consists of the following steps:
@@
    Discard raters with an intra-rater consistency lower than $\mu_{intraL}$ or an inter-rater consistency lower than $\mu_{inter}$ w.r.t. the average scores of the control questions.
Due to the limited number of repeated control questions, small variations in the answers of a rater could have a significant impact on their intra-rater agreement [temel2017samplesize]. Thus, an initial selection based only on high intra-rater consistency could result in discarding too many raters. Using both coefficients, we retain raters who align with the group opinion, even with minor inconsistencies in their responses. This approach improves the overall reliability and representativeness of the data by filtering out erratic and atypical ratings without being excessively restrictive. Figure 2 shows the consistency map of the raters selected following the described procedure, considering values of $0.4$, $0.1$ and $0.2$ for $\mu_{intraH}$, $\mu_{intraL}$ and $\mu_{inter}$, respectively. The responses of 8 raters were discarded due to very low intra-rater consistency. Additionally, 6 raters presented significant deviations from the mean opinion according to $\mu_{inter}$. The resulting set of 22 raters yields a final dataset of 4402 scored trajectories.